# Enumeration of maps with tight boundaries and the Zhukovsky transformation

### v1

@@ 3.1 Adding an extra boundary-vertex
$\begin{split}\frac{\partial A_{L,\ell}}{\partial t}&=\frac{\partial\ }{\partial t}\left([z^{\ell}]\left(z+S+\frac{R}{z}\right)^{L}\right)\\ &=[z^{\ell}]\ L\left(z+S+\frac{R}{z}\right)^{L-1}\left(\frac{\partial S}{\partial t}+\frac{1}{z}\frac{\partial R}{\partial t}\right)\\ &=L\,A_{L-1,\ell}\frac{\partial S}{\partial t}+L\,A_{L-1,\ell+1}\frac{\partial R}{\partial t}.\end{split}$ (51)
@@
$\begin{split}\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t}&=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}\\ &+\sum_{i=1}^{n}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots\frac{\partial A_{L_{i},\ell_{i}}}{\partial t}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{n}}\ .\\ \end{split}$ (55)
@@
$\begin{split}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots&\left(\sum_{m_{i}>\ell_{i}}A_{L_{i},m_{i}}\,m_{i}\,T^{(0)}_{m_{i},0|\ell_{i}}\right)\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{i}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{i},\cdots\ell_{n}}\\ =\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots&A_{L_{i},\ell_{i}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{i}\cdots\ell_{n})\left(\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},0|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots\ell_{n}}\right)\end{split}$ (56)
@@
$\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t}=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}\ ,$ (57)
@@
$T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}=\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}+\sum_{i=1}^{n}\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},0|m_{i}}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}.$ (58)
This identity has the following nice combinatorial interpretation: consider a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}$, and let $v$ be the boundary-vertex corresponding to the last 00. We claim that there exists at most one $i=1,\ldots,n$ such that there exists a closed path of length $m_{i}<\ell_{i}$ separating $v$ and the $i$-th boundary from the others. Indeed, if two such $i_{1}$ and $i_{2}$ existed, we could construct a closed path contradicting the tightness of the $i_{1}$-th or the $i_{2}$-th boundary (see Figure 4). If no such $i$ exists, $v$ can be changed into a regular vertex without affecting the tightness of the other boundaries, and we get a map counted by $\frac{\partial}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. Otherwise, if exactly one such $i$ exists, we can consider the “extremal” shortest closed path separating $v$ and the $i$-th boundary from the others. By cutting along this path, we get on the one hand a map counted by $T^{(0)}_{\ell_{i},0|m_{i}}$ (it is a pair of pants with one tight boundary of length $\ell_{i}$, one boundary-vertex and one strictly tight boundary of length $m_{i}$) and on the other hand a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}$. There are $m_{i}$ ways to glue back these maps together.
Note that the above reasoning also works when $\ell_{i}$ is allowed to take the value 00, in which case the sum over $m_{i}$ vanishes. Putting things together, we arrive at the following:
@@
$\begin{split}T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}&=\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}+\sum_{i=1}^{n}\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},0|m_{i}}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\\ &=\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}+\sum_{i=1}^{n}\bigg{(}\sum_{m_{i}=1\atop m_{i}=\ell_{i}-1\!\!\!\!\mod 2}^{\ell_{i}-1}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}-m_{i}-1}{2}}\frac{\partial S}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\\ &\hskip 93.89418pt+\sum_{m_{i}=1\atop m_{i}=\ell_{i}-2\!\!\!\!\mod 2}^{\ell_{i}-2}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}-m_{i}}{2}-1}\frac{\partial R}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\bigg{)}\,.\\ \end{split}$ (59)
@@ 3.2 Adding an extra boundary-face
$L_{n+1}\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t_{L_{n+1}}}=F^{(g)}_{L_{1},\ldots,L_{n},L_{n+1}}=\sum_{\ell_{1},\ldots,\ell_{n+1}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n+1},\ell_{n+1}}\,(\ell_{1}\cdots\ell_{n+1})\,T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}\ .$ (63)
@@
$\begin{split}L_{n+1}\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t_{L_{n+1}}}&=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})L_{n+1}\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t_{L_{n+1}}}\\ &+\sum_{i=1}^{n}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots\left(L_{n+1}\frac{\partial A_{L_{i},\ell_{i}}}{\partial t_{L_{n+1}}}\right)\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{n}}\ .\\ \end{split}$ (64)
@@
where, to go from the first to the second line, we exchanged the dummy summation variables $\ell_{i}$ and $m_{i}$, as well as the order of summation. As for the first term in the right hand side of (64), using (62), it may be rewritten as
$\sum_{\ell_{1},\ldots,\ell_{n+1}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n+1},\ell_{n+1}}(\ell_{1}\cdots\ell_{n+1})D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}\ .$ (66)
@@
$T^{(g)}_{\ell_{1},\ldots,\ell_{n},\ell_{n+1}}=D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}+\sum_{i=1}^{n}\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},\ell_{n+1}|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots\ell_{n}}\ .$ (68)
@@
$\begin{split}T^{(g)}_{\ell_{1},\ldots,\ell_{n},\ell_{n+1}}&=D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}+\sum_{i=1}^{n}\bigg{(}\sum_{m_{i}=1\atop m_{i}=\ell_{i}-1\!\!\!\!\mod 2}^{\ell_{i}-1}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}-1}{2}}\frac{\partial S}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\\ &\hskip 93.89418pt+\sum_{m_{i}=1\atop m_{i}=\ell_{i}-2\!\!\!\!\mod 2}^{\ell_{i}-2}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}-1}\frac{\partial R}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\bigg{)}\,,\\ \end{split}$ (69)
@@
$\begin{split}T^{(g)}_{\ell_{1},\ldots,\ell_{n},\ell_{n+1}}&=D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}+\sum_{i=1}^{n}\bigg{(}\sum_{m_{i}=1\atop m_{i}=\ell_{i}-1\!\!\!\!\mod 2}^{\ell_{i}-1}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}-1}\frac{\partial R}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\\ &\hskip 93.89418pt+\sum_{m_{i}=1\atop m_{i}=\ell_{i}-2\!\!\!\!\mod 2}^{\ell_{i}-2}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}-1}{2}}\frac{\partial S}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\bigg{)}\,.\\ \end{split}$ (70)
@@
Again, the identity (68) has a clear combinatorial interpretation. Consider a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}$. Then, there exists at most one $i\in\{1,\ldots,n\}$ such that there exists a closed path of length $m_{i}<\ell_{i}$ separating the $i$-th and the $(n+1)$-th boundaries from the others, for a reason similar to that illustrated in Figure 4. If exactly one such $i$ exists, we can consider the extremal shortest separating closed path and cut along it. We get, on the one hand, a map counted by $T^{(0)}_{\ell_{i},\ell_{n+1}|m_{i}}$, i.e. a pair of pants with two tight boundaries of lengths $\ell_{i}$ and $\ell_{n+1}$, and one strictly tight boundary of length $m_{i}$, and, on the other hand, a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}$. There are $m_{i}$ ways to glue back these maps together. If no such $i$ exists, the $(n+1)$-th boundary can be changed into an inner face without affecting the tightness of the other boundaries. We deduce that such maps are counted by $D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$, which explains, in retrospect, the name of the operator $D_{m}$.
@@
For $g=0$ and $n=2$, notice that Proposition 3.5 relates the generating function of tight cylinders, given by $T^{(0)}_{\ell_{1},\ell_{2}}=\delta_{\ell_{1},\ell_{2}}R^{\ell_{1}}/\ell_{1}$ for $\ell_{1}>0$, and $T^{(0)}_{0,0}=\ln(R/t)$, with that of tight pairs of pants given by Theorem A.1 below.
@@ 3.3 Consistency with the expression for the planar bipartite case
In this section we check that the explicit expression (32) for $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ in the planar bipartite case is consistent with the above relations (59) and (68) describing the addition of an extra boundary-vertex and an extra boundary-face respectively.
@@
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},0}\right|_{\mathrm{bip}}=\frac{\partial}{\partial t}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ (71)
@@
which is easily derived from the combinatorial interpretation of the Bell polynomials in terms of partitions, we deduce that
@@
$\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}-1}R^{\prime}\times\\ \sum_{k=0}^{n-2}\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})k!p_{k}(\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n})b_{n-2,k+1}.$ (76)
@@
$(k+1)p_{k+1}(\ell_{1},\ldots,\ell_{n})=(\ell_{1}+\cdots+\ell_{n}-k-1)p_{k}(\ell_{1},\ldots,\ell_{n})\\ +\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})p_{k}(\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n})\ .$ (77)
This latter recursion is proved in [polytightmaps, Proposition 2.2].
@@
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=D_{2\ell_{n+1}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}+\ell_{n+1}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}.$ (78)
@@
Here, we may justify the first equality by noting that, in the bipartite case, $R^{(j)}$ can be seen as the generating function of maps with one unrooted boundary-face of length $2$ and $j+1$ boundary-vertices. Since $D_{2m}$ creates an extra tight boundary-face of length $2m$, and since the other boundaries (which have lengths $2$ and 00 only) are automatically tight, we indeed obtain the maps counted by $T^{(0)}_{2m,2,\!\!\!\!\underbrace{\scriptstyle 0,\ldots,0}_{j+1\text{ times}}}\Big{|}_{\mathrm{bip}}$. As for the second equality, it is a particular instance of (32), simplified via the identity $p_{k}(m,1,0,\ldots,0)=p_{k}(m,1)=p_{k}(m)+p_{k-1}(m)=q_{k}(m)$.
@@
(throughtout this computation, we will leave the summation ranges as implicit as possible, since they are naturally enforced by the vanishing of the summands). On the other hand, we have, using (79),
$D_{2\ell_{n+1}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}-1}\sum_{k\geq 0}k!p_{k}(\ell_{1},\ldots,\ell_{n})\times\\ \left((\ell_{1}+\cdots+\ell_{n}-k-1)(D_{2\ell_{n+1}}R)b_{n-2,k+1}+\sum_{j\geq 1}(D_{2\ell_{n+1}}R^{(j)})b_{n-2,k+1}^{[j]}\right)\\ =R^{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1}}\sum_{k\geq 0}k!p_{k}(\ell_{1},\ldots,\ell_{n})\times\hfill\\ \left((\ell_{1}+\cdots+\ell_{n}-k-1)\frac{R^{\prime}}{R}b_{n-2,k+1}+\sum_{j\geq 1}\sum_{k^{\prime}\geq 0}k^{\prime}!q_{k^{\prime}}(\ell_{n+1})b_{j+1,k^{\prime}+1}b_{n-2,k+1}^{[j]}\right)\ .$ (81)
@@
upon interpreting the lhs as a sum over partitions of $\{1,\ldots,n-1\}$ into $k^{\prime}+k^{\prime\prime}+1$ blocks, $k^{\prime}+1$ of which are marked, the element $n-1$ being always in a marked block: the first equality is obtained by summing over the number $j$ of elements other than $n-1$ that are in a marked block, and the second equality is obtained by noting that $\binom{n-2}{j}b_{n-2-j,k^{\prime\prime}}=b_{n-2,k^{\prime\prime}+1}^{[j]}$ for $j\geq 1$. We deduce that the sum appearing at the end of (81), with $k$ replaced by $k^{\prime\prime}$, can be rewritten as
@@
We may similarly check that the explicit expression (44) for $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ in the planar quasi-bipartite case is also consistent with the relations (59) and (68) for the addition of an extra boundary. This property is now a direct consequence of the string equation relating $\tilde{p}_{k+1}(\ell_{1},\ell_{2};\ell_{3},\ldots,\ell_{n})$ to $\tilde{p}_{k}(\ell_{1},\ell_{2};\ell_{3},\ldots,\ell_{n})$, see [polytightmaps, Proposition 2.10].

### v2

@@ 3.1 Adding an extra boundary-vertex
$\frac{\partial A_{L,\ell}}{\partial t}=\frac{\partial\ }{\partial t}\left([z^{\ell}]\left(z+S+\frac{R}{z}\right)^{L}\right)\\ =[z^{\ell}]\ L\left(z+S+\frac{R}{z}\right)^{L-1}\left(\frac{\partial S}{\partial t}+\frac{1}{z}\frac{\partial R}{\partial t}\right)=L\,A_{L-1,\ell}\frac{\partial S}{\partial t}+L\,A_{L-1,\ell+1}\frac{\partial R}{\partial t}.$ (52)
@@
[…]
@@
$\begin{split}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots&\left(\sum_{m_{i}>\ell_{i}}A_{L_{i},m_{i}}\,m_{i}\,T^{(0)}_{m_{i},0|\ell_{i}}\right)\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{i}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{i},\ldots\ell_{n}}\\ =\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots&A_{L_{i},\ell_{i}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{i}\cdots\ell_{n})\left(\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},0|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots\ell_{n}}\right)\end{split}$ (57)
@@
$\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t}=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}\$, (58)
@@
$T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}=\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}+\sum_{i=1}^{n}\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},0|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots,\ell_{n}}$. (59)
This identity has the following nice combinatorial interpretation: consider a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}$, and let $v$ be the boundary-vertex corresponding to the last 0. We claim that there exists at most one $i=1,\ldots,n$ such that there exists a closed path of length $m_{i}<\ell_{i}$ separating $v$ and the $i$-th boundary from the others. Indeed, if two such $i_{1}$ and $i_{2}$ existed, we could construct a closed path contradicting the tightness of the $i_{1}$-th or the $i_{2}$-th boundary (see Figure 4). If no such $i$ exists, $v$ can be changed into a regular vertex without affecting the tightness of the other boundaries, and we get a map counted by $\frac{\partial}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. Otherwise, if exactly one such $i$ exists, we can consider the “extremal” shortest closed path separating $v$ and the $i$-th boundary from the others. By cutting along this path, we get on the one hand a map counted by $T^{(0)}_{\ell_{i},0|m_{i}}$ (it is a pair of pants with one tight boundary of length $\ell_{i}$, one boundary-vertex and one strictly tight boundary of length $m_{i}$) and on the other hand a map counted by $T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots,\ell_{n}}$. There are $m_{i}$ ways to glue these maps back together.
Note that the above reasoning also works when $\ell_{i}$ is allowed to take the value 0, in which case the sum over $m_{i}$ vanishes. Putting things together, we arrive at the following proposition.
@@
[…]
@@ 3.2 Adding an extra boundary-face
$L_{n+1}\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t_{L_{n+1}}}=F^{(g)}_{L_{1},\ldots,L_{n},L_{n+1}}=\sum_{\ell_{1},\ldots,\ell_{n+1}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n+1},\ell_{n+1}}\,(\ell_{1}\cdots\ell_{n+1})\,T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}\$. (64)
@@
[…]
@@
where, to go from the first to the second line, we exchanged the dummy summation variables $\ell_{i}$ and $m_{i}$, as well as the order of summation. As for the first term on the right hand side of (65), using (63), it may be rewritten as
$\sum_{\ell_{1},\ldots,\ell_{n+1}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n+1},\ell_{n+1}}(\ell_{1}\cdots\ell_{n+1})D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}\$. (67)
@@
$T^{(g)}_{\ell_{1},\ldots,\ell_{n},\ell_{n+1}}=D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}+\sum_{i=1}^{n}\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},\ell_{n+1}|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots\ell_{n}}\$. (69)
@@
[…]
@@
[…]
@@
Again, the identity (69) has a clear combinatorial interpretation. Consider a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}$. Then, there exists at most one $i\in\{1,\ldots,n\}$ such that there exists a closed path of length $m_{i}<\ell_{i}$ separating the $i$-th and the $(n+1)$-th boundaries from the others, for a reason similar to that illustrated in Figure 4. If exactly one such $i$ exists, we can consider the extremal shortest separating closed path and cut along it. We get, on the one hand, a map counted by $T^{(0)}_{\ell_{i},\ell_{n+1}|m_{i}}$, i.e. a pair of pants with two tight boundaries of lengths $\ell_{i}$ and $\ell_{n+1}$, and one strictly tight boundary of length $m_{i}$, and, on the other hand, a map counted by $T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots,\ell_{n}}$. There are $m_{i}$ ways to glue back these maps together. If no such $i$ exists, the $(n+1)$-th boundary can be changed into an inner face without affecting the tightness of the other boundaries. We deduce that such maps are counted by $D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$, which explains, in retrospect, the name of the operator $D_{m}$.
@@
For $g=0$ and $n=2$, notice that Proposition 3.5 relates the generating function of tight cylinders $T^{(0)}_{\ell_{1},\ell_{2}}$, given by (25) for $\ell_{1},\ell_{2}>0$, and $T^{(0)}_{0,0}=\ln(R/t)$, with that of tight pairs of pants given by Theorem A.1 below.
@@ 3.3 Consistency with the expression for the planar bipartite case
In this section we take $n\geq 3$ and check that the explicit expression (33) for $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ in the planar bipartite case is consistent with the above relations (60) and (69) describing the addition of an extra boundary-vertex and an extra boundary-face respectively.
@@
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},0}\right|_{\mathrm{bip}}=\frac{\partial}{\partial t}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2m_{i},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ (72)
@@
which is easily derived from the combinatorial interpretation of the Bell polynomials in terms of partitions (with the convention that $b_{n,k}=0$ if $k>n$), we deduce that
@@
$\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2m_{i},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}-1}R^{\prime}\times\\ \sum_{k=0}^{n-2}\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})k!p_{k}(\ell_{1},\ldots,m_{i},\ldots,\ell_{n})b_{n-2,k+1}.$ (77)
@@
$(k+1)p_{k+1}(\ell_{1},\ldots,\ell_{n})=(\ell_{1}+\cdots+\ell_{n}-k-1)p_{k}(\ell_{1},\ldots,\ell_{n})\\ +\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})p_{k}(\ell_{1},\ldots,m_{i},\ldots,\ell_{n})\ .$ (78)
This latter recursion is proved in [BGM24, Proposition 2.2].
@@
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=D_{2\ell_{n+1}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}+\ell_{n+1}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2m_{i},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}.$ (79)
@@
Here, we may justify the first equality by noting that, in the bipartite case, $R^{(j)}$ can be seen as the generating function of maps with one unrooted boundary-face of length $2$ and $j+1$ boundary-vertices. Since $D_{2m}$ creates an extra tight boundary-face of length $2m$, and since the other boundaries (which have lengths $2$ and 0 only) are automatically tight, we indeed obtain the maps counted by $T^{(0)}_{2m,2,\!\!\!\!\underbrace{\scriptstyle 0,\ldots,0}_{j+1\text{ times}}}\Big|_{\mathrm{bip}}$. As for the second equality, it is a particular instance of (33), simplified via the identity $p_{k}(m,1,0,\ldots,0)=p_{k}(m,1)=p_{k}(m)+p_{k-1}(m)=q_{k}(m)$.
@@
(throughtout this computation, we will leave the summation ranges as implicit as possible, since they are naturally enforced by the vanishing of the summands). On the other hand, using (80), we have
$D_{2\ell_{n+1}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}-1}\sum_{k\geq 0}k!p_{k}(\ell_{1},\ldots,\ell_{n})\times\\ \allowdisplaybreaks\left((\ell_{1}+\cdots+\ell_{n}-k-1)(D_{2\ell_{n+1}}R)b_{n-2,k+1}+\sum_{j\geq 1}(D_{2\ell_{n+1}}R^{(j)})b_{n-2,k+1}^{[j]}\right)\\ =R^{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1}}\sum_{k\geq 0}k!p_{k}(\ell_{1},\ldots,\ell_{n})\times\hfill\\ \left((\ell_{1}+\cdots+\ell_{n}-k-1)\frac{R^{\prime}}{R}b_{n-2,k+1}+\sum_{j\geq 1}\sum_{k^{\prime}\geq 0}k^{\prime}!q_{k^{\prime}}(\ell_{n+1})b_{j+1,k^{\prime}+1}b_{n-2,k+1}^{[j]}\right)\ .$ (82)
@@
upon interpreting the lhs as a sum over partitions of $\{1,\ldots,n-1\}$ into $k^{\prime}+k^{\prime\prime}+1$ blocks, $k^{\prime}+1$ of which are marked, the element n1n-1 being always in a marked block: the first equality is obtained by summing over the number $j$ of elements other than n1n-1 that are in a marked block, and the second equality is obtained by noting that $\binom{n-2}{j}b_{n-2-j,k^{\prime\prime}}=b_{n-2,k^{\prime\prime}+1}^{[j]}$ for $j\geq 1$. We deduce that the sum appearing at the end of (82), with $k$ replaced by $k^{\prime\prime}$, can be rewritten as
@@
We may similarly check that the explicit expression (45) for $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ in the planar quasi-bipartite case is also consistent with the relations (60) and (69) for the addition of an extra boundary. This property is now a direct consequence of the string equation relating $\tilde{p}_{k+1}(\ell_{1},\ell_{2};\ell_{3},\ldots,\ell_{n})$ to $\tilde{p}_{k}(\ell_{1},\ell_{2};\ell_{3},\ldots,\ell_{n})$, see [BGM24, Proposition 2.10].